### v1

# Finding Kissing Numbers with Game-theoretic Reinforcement Learning

Abstract
Since Isaac Newton first studied the Kissing Number Problem in 1694, determining the maximal number of non-overlapping spheres around a central sphere has remained a fundamental challenge. This problem represents the local analogue of Hilbert’s 18th problem on sphere packing, bridging geometry, number theory, and information theory. Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry and exponentially growing combinatorial complexity beyond 8 dimensions, which exceeds the complexity of Go game, limit the scalability of existing methods. Here we model this problem as a two-player matrix completion game and train the game-theoretic reinforcement learning system, PackingStar, to efficiently explore high-dimensional spaces. The matrix entries represent pairwise cosines of sphere center vectors; one player fills entries while another corrects suboptimal ones, jointly maximizing the matrix size, corresponding to the kissing number. This cooperative dynamics substantially improves sample quality, making the extremely large spaces tractable. PackingStar reproduces previous configurations and surpasses all human-known records from dimensions 25 to 31, with the configuration in 25 dimensions geometrically corresponding to the Leech lattice and suggesting possible optimality. It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions and discovers over 6000 new structures in 14 and other dimensions. These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition and open new pathways for the Kissing Number Problem and broader geometry problems.
Introduction
The Kissing Number Problem was first posed in 1694 by Isaac Newton and David Gregory, asking for the maximal number of non-overlapping spheres that can touch a central sphere in three-dimensional space. It was later extended to $n$-dimensional Euclidean space, where $K(n)$ denotes the kissing number in $n$ dimensions, sparking centuries of research in high-dimensional geometry. In 1900, David Hilbert included the sphere packing problem in $n$-dimensional space as the 18th problem of his famous 23 problems [?]. The Kissing Number Problem can be viewed as a local analogue of Hilbert’s 18th problem [?]. Beyond geometry, it connects to number theory, group theory, and information theory [?,?,?]. Therefore, advancing this problem could foster progress across multiple fields.
Table 1: Kissing number $K(n)$ in dimension $n$ for $1\leq n\leq 24$ [?]. The bold items indicate the kissing numbers that are known to be optimal.
$n$ $K(n)$ $n$ $K(n)$ $n$ $K(n)$ $n$ $K(n)$ $n$ $K(n)$ $n$ $K(n)$
1 2 5 40 9 306 13 1154 17 5730 21 29768
2 6 6 72 10 510 14 1932 18 7654 22 49896
3 12 7 126 11 593 15 2564 19 11692 23 93150
4 24 8 240 12 840 16 4320 20 19448 24 196560
Unfortunately, determining kissing numbers in higher dimensions remains exceptionally difficult. Table 1 lists the kissing numbers in different dimensions. After centuries of effort, the problem has been solved exactly only in a few dimensions 1, 2, 3, 4, 8, and 24 [?,?,?,?,?,?,?]. As the dimension increases, human understanding becomes more limited due to the growing geometric complexity of high-dimensional spaces and combinatorial explosion of configurations. Over the past half-century, mathematicians have continued to push the boundaries using lattice constructions [?], code techniques [?], group representation theory [?], and carefully designed substructures derived from known configurations [?]. In dimensions 10, 11, 13, 14, and dimensions from 17 to 21, such approaches established new lower bounds [?,?]. In dimensions 25–31, studies conducted in 2011 and 2016 further improved the bounds by exploiting subsets of Leech lattice vectors combined with heuristic optimization [?,?]. While these results mark significant milestones, they are limited by reliance on prior knowledge, and painstaking human design. As dimensionality increases, the exponentially growing configuration space and the breakdown of full symmetry assumptions make existing approaches insufficient. These challenges highlight the need for novel paradigms to explore complex structures beyond human intuition.
In this paper, we introduce PackingStar, a multi-agent reinforcement learning system that reformulates the Kissing Number Problem as a two-player matrix completion game. This approach scales efficiently across dimensions, mitigates the combinatorial explosion of high-dimensional search spaces, and overcomes the challenge of extreme data scarcity due to the limited human knowledge of high-dimensional constructions. In PackingStar, two cooperative agents learn to play the matrix completion game to efficiently explore irregular structures in high-dimensional spaces without human prior knowledge. The first agent focuses on incrementally filling matrix entries to construct larger and denser configurations. However, in the early stages of exploration, the candidate space for potential entries is exponentially large, and the agent only observes local matrix information, making erroneous placements inevitable. To address this problem, the second agent is introduced to learn how to identify and remove invalid or inconsistent entries. By leveraging a more global understanding of the matrix geometry and high-dimensional structure, this agent dynamically corrects the mistakes made by the first agent, enabling the joint system to converge toward larger configurations—corresponding to higher kissing numbers. The key insight to PackingStar’s success lies in two complementary design principles. First, the framework reformulates the kissing number construction as an equivalent matrix completion problem in the Gram matrix domain. This reformulation eliminates numerical instability and inefficiency inherent in high-dimensional coordinate representations, enabling all computations to operate directly on the Gram matrix in a fully parallelizable manner. The formulation seamlessly couples problem modeling with its learning methods, allowing efficient GPU acceleration and large-scale sampling for AI-driven discovery of kissing numbers. Second, PackingStar features the finely designed game-theoretic interaction between the two agents. By strategically performing delayed and selective corrections, the second agent imposes geometric constraints on the first agent’s expansive search, effectively regularizing exploration in an otherwise intractable space. This game dynamics not only improves the quality of sampled configurations and the efficiency of value estimation but also reduces the search complexity in exponentially large spaces, enabling scalable discovery of complex high-dimensional configurations that were previously inaccessible to existing methods.
Figure 1: AI breakthroughs in the Kissing Number Problem. Grey markers show the best previously known lower bounds. The blue triangle marks the 11D improvement achieved by AlphaEvolve (2025) [?]. In 13D, PackingStar (red star) uncovers a new rational configuration with kissing number $K_{r}(13)=1146$, distinct from the current best non-rational lower bound $K(13)=1154$. For $25\leq n\leq 31$, the red stars indicate new lower bounds $K_{\mathrm{new}}(n)$ discovered by PackingStar, with the inset reporting the corresponding increments $\Delta K(n)=K_{\mathrm{new}}(n)-K_{\mathrm{prev}}(n)$ compared to the previous best $K_{\mathrm{prev}}(n)$.
PackingStar demonstrates remarkable scalability, successfully reproducing all known kissing number configurations in dimensions up to 31. Figure 1 illustrates some of PackingStar’s breakthrough discoveries on this problem. From dimensions 25 to 31, PackingStar surpasses the lower bounds reported in 2016 and also exceeds the meta-construction form introduced in 2011 [?,?], achieving what appears to be the optimal construction form along this path; the resulting construction form are summarized in Table 2. Notably, the breakthrough configuration discovered in 25 dimensions exhibits a clear geometric pattern that corresponds closely to the substructure of Leech lattice. Across multiple metrics, this configuration shows strong convergence, suggesting that it may represent the optimal 25-dimensional arrangement, although a rigorous mathematical proof is still lacking. In 13 dimensions, PackingStar surpasses the previously known rational structure with 1130 spheres from 1971 [?], reaching a kissing number of 1146. Rational structures are those in which all pairwise cosines between sphere centers are rational. The non-rational structure with 1154 spheres is obtained by slightly adjusting the coordinates of some spheres in the 1130-sphere configuration, later reached 1154 in 1999 [?]. Although the new 13-dimensional rational configuration does not exceed the non-rational record, it is of considerable mathematical value. Its fully rational inter-sphere cosines allow for exact analytic study without relying on numerical approximations. This offers insights into the effects of different constraints on kissing numbers and may inspire higher-density packings. Rational structures also facilitate more efficient encoding and decoding in information and communication applications, highlighting their practical significance alongside their theoretical and geometric value. Beyond these specific dimensions, PackingStar uncovers over 6000 new configurations in 14 dimensions and identifies additional configurations corresponding to current lower bounds in dimensions 12, 13, 15, and others. The sheer volume of these newly discovered arrangements provides rich opportunities for further investigation and may ultimately lead to configurations with even higher kissing numbers. Collectively, these results highlight AI’s ability to understand and reason about high-dimensional geometric structures, opening a new paradigm for studying the Kissing Number Problem and demonstrating the potential of AI to explore a broader class of challenging mathematical problems.
Key difference to AlphaEvolve
The key difference between PackingStar and AlphaEvolve lies in the search space. AlphaEvolve optimizes explicit coordinates in $\mathbb{R}^{n}$, which is numerically unstable and scales exponentially with dimension; increasing computational resources cannot overcome this fundamental limitation. In contrast, PackingStar operates in the Gram-matrix domain with strict geometric constraints, enabling efficient large-scale sampling. This fundamental distinction explains why PackingStar scales to higher dimensions while AlphaEvolve cannot [?].
Although AlphaEvolve improved the kissing number in 11 dimensions from 592 to 593, likely by adjusting the coordinates of the known 592-sphere configuration, the result lacks clear geometric structure and does not generalize to other dimensions. In contrast, PackingStar not only sets new bounds across higher dimensions but also discovers geometrically meaningful, analyzable configurations, providing concrete objects for further mathematical study [?].
Results
We evaluate the PackingStar system across multiple high-dimensional regimes and report new state-of-the-art kissing configurations. Our results fall into three categories: (i) new lower bounds in dimensions 25253131, (ii) improved rational constructions in $13$ dimensions, and (iii) diverse families of solutions in different dimensions. These findings demonstrate that PackingStar can both recover classical lattice-based structures and uncover entirely new high-dimensional arrangements, substantially expanding the landscape of known kissing configurations. The main kissing configurations are available at the data repository https://github.com/CDM1619/PackingStar.
New records from dimensions 25 to 31
We first evaluate PackingStar in dimensions $25\leq n\leq 31$, where the best previously known lower bounds were all derived from a common structural template that the Leech lattice packing together with a family of disjoint subsets $S_{i}$ of its minimal vectors [?,?].
Table 2: New kissing numbers $K_{\mathrm{new}}(n)$ by PackingStar vs. previous $K_{\mathrm{prev}}(n)$ for $25\leq n\leq 31$. The asterisk $\bigstar$ indicates that under our new construction form, the theoretical lower bound in 31 dimensions should be 238350. The present version reports 238078 only due to computational and time limitations, rather than any intrinsic limitation of our method. The Future versions are expected to reach the theoretical bound. In the new construction form, bold items indicate the improved configuration forms discovered by PackingStar compared with previous forms.
   $n$    New construction form    $\boldsymbol{K_{\mathrm{new}}(n)}$    $K_{\mathrm{prev}}(n)$ [?]
   25    $K(24)+|S_{1}|$    197056    197048
   26    $\boldsymbol{K(2)}+K(24)+2|S_{1}|+2|S_{2}|$    198550    198512
   27    $\boldsymbol{K(3)}+K(24)+2|S_{1}|+2|S_{2}|+\sum_{i=3}^{5}|S_{i}|$    200044    199976
   28    $\boldsymbol{K(4)}+K(24)+2\sum_{i=1}^{8}|S_{i}|$    204520    204368
   29    $\boldsymbol{K(5)}+K(24)+\boldsymbol{2\sum_{i=1}^{12}|S_{i}|+\sum_{i=13}^{14}|S_{i}|}$    209496    208272
   30    $\boldsymbol{K(6)}+K(24)+2\sum_{i=1}^{24}|S_{i}|$    220440    219984
   31    $\boldsymbol{K(7)}+K(24)+\boldsymbol{2\sum_{i=1}^{42}|S_{i}|}$       238078    232874
In this setting, the first objective is to identify the largest possible subsets $S_{i}$. Each $S_{i}$ corresponds to a subset of shortest Leech lattice vectors whose pairwise cosine values are bounded by $1/4$; equivalently, in the PackingStar framework this corresponds to cosine sets $C_{1}=C_{2}=\{-1,\ 0,\ \pm 1/4\}$ together with the structural constraint that $C_{*}$ matches the Leech minimal-vector set. Previous work identified subsets $|S_{i}|=488$ with no apparent geometric structure [?]. PackingStar discovers subsets $|S_{i}|=496$ exhibiting a clear internal geometric organization. Specifically, this 496-element subset is composed of 28 symmetric-frame structures $\mathrm{X}_{8}$ in 8-dimensional subspaces, each containing 16 spheres, together with a symmetric-frame structure (“Conway-Curtis Cross” [?]) $\mathrm{X}_{24}$ in 24-dimensional space containing 48 spheres, which corresponds to an important substructure of the Leech lattice. Here the symmetric-frame structure is a Cross-polytope. Based on this clear geometric interpretation, we conjecture that this subset $S_{i}$ may be optimal, and consequently, the 25-dimensional kissing configuration derived from this construction could also attain the maximum kissing number.
The second objective is to identify improved meta construction forms that assemble these subsets into full configurations. Both objectives are jointly addressed within the unified PackingStar framework, which simultaneously increases the achievable size of $S_{i}$ and uncovers superior meta construction forms. Our findings show that PackingStar not only discovers larger subsets but also identifies a strictly better construction template than the meta-construction introduced in 2011; moreover, evidence from convergence and geometric consistency strongly suggests that the new template is optimal within this construction pathway. These advances yield especially pronounced improvements in dimensions 29 and 31. Table 2 reports additional new construction forms that deviate from the classical form [?,?]. In dimensions 26263030, the system consistently augments the Leech-lattice component by adding lower-dimensional parts $K(k)$ where k=n24k=n-24 (e.g., adding $K(2)$ in dimension $26$). In 31 dimensions, PackingStar uncovers a novel assembly pattern by partitioning a 7-dimensional kissing configuration into 42 (optimal) disjoint unit-radius equilateral triangles, resulting in an $84$-fold weighted $S_{i}$, surpassing the previous $75$-fold weighted $S_{i}$. Similarly, in 29 dimensions, PackingStar embeds 12 (optimal) disjoint unit-radius equilateral triangles into the 5-dimensional kissing configuration to generate a $26$-fold weighted $S_{i}$, exceeding the prior $24$-fold weighted $S_{i}$. These advances yield strictly larger kissing configurations in all dimensions from $25$ to $31$. The resulting new lower bounds $K_{\mathrm{new}}(n)$ are summarized in Table 2. Compared with the previous best values $K_{\mathrm{prev}}(n)$, our constructions increase the kissing number by $8,38,68,152,1224,456,$ and $5204$ spheres in dimensions $25$ through $31$, respectively. These improvements are also highlighted by the red markers in Fig. 1. Table 3 summarizes the cosine value among sphere pairs in the new configurations.
Improved rational configurations in 13 dimensions
We apply PackingStar to explore rational kissing configurations in dimension $13$. In this setting, guided by the simulations in Step 1 of our method, we identify two feasible cosine sets for 13-dimensional rational arrangements: $C_{1}=C_{2}=\{-1,\ 0,\ \pm 1/4,\ \pm 1/2\}$ or $\{-1,\ -3/4,\ 0,\ \pm 1/4,\ \pm 1/2\}$. Table 3 summarizes the cosine value among sphere pairs in $K_{r}(13)$ configurations. Based on these cosine features, PackingStar discovers several new rational configuration with kissing number $K_{r}(13)=1146$, surpassing the previously known rational structure from 1971, which achieved $1130$ [?]. The new configuration exhibits a highly regular geometric pattern and a fully rational Gram matrix, allowing exact verification of all pairwise cosines without reliance on numerical approximations. This demonstrates the capability of PackingStar to systematically explore the space of rational configurations and discover previously unknown structures.
Table 3: Cosine sets of different kissing configurations. For each configuration, the table lists the set of cosine values between distinct unit vectors. All maximal cosines are $\leq 0.5$, ensuring that the geometric constraints of the kissing number problem are satisfied, i.e., no spheres overlap.
   Configuration    Cosine set (excluding 1-1 and 0)
   $K_{\mathrm{r}}(13)$    $\{\pm\tfrac{1}{4},\ \pm\tfrac{1}{2}\}$; variants: $\{-\tfrac{3}{4},\ \pm\tfrac{1}{4},\ \pm\tfrac{1}{2}\}$
   $K_{\mathrm{new}}(25)$    $\{\pm\tfrac{1}{6},\ \pm\tfrac{\sqrt{6}}{12},\ \pm\tfrac{1}{4},\ \pm\tfrac{1}{3},\ \pm\tfrac{\sqrt{6}}{6},\ \pm\tfrac{1}{2}\}$
   $K_{\mathrm{new}}(26)$    $\{-\tfrac{5}{6},\ -\tfrac{2}{3},\ \pm\tfrac{1}{6},\ \pm\tfrac{\sqrt{6}}{12},\ \pm\tfrac{1}{4},\ \pm\tfrac{1}{3},\ \pm\tfrac{\sqrt{6}}{6},\ \pm\tfrac{1}{2}\}$
   $K_{\mathrm{new}}(27)$    $\{-\tfrac{5}{6},\ -\tfrac{2}{3},\ \pm\tfrac{2\sqrt{3}-\sqrt{6}}{12},\ \pm\tfrac{1}{6},\ \pm\tfrac{\sqrt{6}}{12},\ \pm\tfrac{1}{4},\ \pm\tfrac{1}{3},\ \pm\tfrac{\sqrt{6}}{6},\ \pm\tfrac{2\sqrt{3}+\sqrt{6}}{12},\ \pm\tfrac{1}{2}\}$
   $K_{\mathrm{new}}(28)$    $\{-\tfrac{5}{6},\ -\tfrac{2}{3},\ \pm\tfrac{1}{6},\ \pm\tfrac{\sqrt{6}}{12},\ \pm\tfrac{1}{4},\ \pm\tfrac{1}{3},\ \pm\tfrac{\sqrt{6}}{6},\ \pm\tfrac{1}{2}\}$
   $K_{\mathrm{new}}(29)$    $\{-\tfrac{5}{6},\ -\tfrac{2}{3},\ \pm\tfrac{2\sqrt{3}-\sqrt{6}}{12},\ \pm\tfrac{1}{6},\ \pm\tfrac{\sqrt{6}}{12},\ \pm\tfrac{1}{4},\ \pm\tfrac{\sqrt{3}}{6},\ \pm\tfrac{1}{3},\ \pm\tfrac{\sqrt{6}}{6},\ \pm\tfrac{2\sqrt{3}+\sqrt{6}}{12},\ \pm\tfrac{1}{2}\}$
   $K_{\mathrm{new}}(30)$    $\{-\tfrac{5}{6},\ -\tfrac{2}{3},\ \pm\tfrac{1}{6},\ \pm\tfrac{\sqrt{6}}{12},\ \pm\tfrac{1}{4},\ \pm\tfrac{1}{3},\ \pm\tfrac{\sqrt{6}}{6},\ \pm\tfrac{1}{2}\}$
   $K_{\mathrm{new}}(31)$    The full cosine set is available at the data repository.
Remarkably, among the 1146-sphere configurations, we observe a variant that can be partitioned into two components: one containing $1008$ spheres and the other $138$ spheres. The $1008$-sphere component exhibits two fundamentally distinct forms: one fully antipodal and the other completely non-antipodal. In the non-antipodal form, none of the spheres has an antipodal counterpart, corresponding to a cosine set of $\{-3/4,0,\pm 1/4,\pm 1/2\}$, excluding 1-1. At the same time, it exhibits high uniformity, with each sphere tightly contacting 98 neighboring spheres. Such a structure is extremely rare and has no known analogues among previously discovered kissing configurations. This reveals that even in 13 and higher dimensions, kissing configurations may possess special non-symmetric arrangements that were previously unexplored.
Diverse constructions in different dimensions
Beyond the specific breakthroughs in 13 and in dimensions 25 through 31, PackingStar uncovers a rich diversity of kissing configurations whose kissing number reaches current lower bound across multiple other dimensions. Among the thousands of discovered variants, we consistently observe the presence of asymmetric structures with cosine set $C_{1}=C_{2}=\{-1,\ -3/4,\ 0,\ \pm 1/4,\ \pm 1/2\}$, similar to those identified in the 13-dimensional rational configuration.
Methods
In this section, we present a learning-based system PackingStar to find high-dimensional kissing constructions. First, we extract dominant geometric features of candidate kissing configurations via simulation and Monte Carlo Tree Search. Second, we formulate the search for maximal kissing numbers as a cooperative two-player matrix completion game, where one agent iteratively fills the Gram matrix and the other corrects suboptimal entries, jointly maximizing the final matrix size.
Identifying geometric features of the kissing configuration via simulation
In this step, we will identify the basic geometric features of possible kissing structures through extensive simulations. This process corresponds to Step 1 in Figure 2. Given $m$ existing spheres in $\mathbb{R}^{n}$ with $m\geq n-1$, we iteratively generate new sphere centers by considering every (n1)(n-1)-combination of existing spheres. Let $\mathbf{x}_{1},\dots,\mathbf{x}_{n-1}\in\mathbb{R}^{n}$ be the corresponding coordinates, and construct $\mathbf{A}=[\mathbf{x}_{1}^{T};\dots;\mathbf{x}_{n-1}^{T}]\in\mathbb{R}^{(n-1)\times n}$, which must have full row rank. We solve for a new sphere center $\mathbf{x}\in\mathbb{R}^{n}$ on the unit sphere satisfying
$\begin{cases}\mathbf{A}\mathbf{x}=\mathbf{b},\\ \|\mathbf{x}\|_{2}=1\end{cases}\quad\mathbf{b}=\tfrac{1}{2}\mathbf{1}_{n-1}.$ (1)
The feasible solutions are
$\mathbf{x}=\mathbf{A}^{+}\mathbf{b}\pm\sqrt{1-\|\mathbf{A}^{+}\mathbf{b}\|_{2}^{2}}\,\mathbf{z},\quad\mathbf{z}\in\ker(\mathbf{A}),\ \|\mathbf{z}\|_{2}=1.$ (2)
All feasible candidates $\mathbf{x}$ are added to the search space, and one is selected as a new center using Monte Carlo Tree Search (MCTS). Each node in the search tree corresponds to a partial configuration $s_{t}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{m_{t}}\}\subset\mathbb{R}^{n\times m_{t}}$ of $m_{t}$ existing sphere centers at step $t$. The MCTS algorithm selects the next candidate $\mathbf{x}_{t}$ according to the upper confidence bound:
$\mathbf{x}_{t}=\arg\max_{\mathbf{x}}\Big[Q(s_{t},\mathbf{x})+c\sqrt{\frac{\ln N(s_{t})}{N(s_{t},\mathbf{x})}}\Big],$ (3)
where $Q(s_{t},\mathbf{x})$ is the estimated value, $N(s_{t})$ and $N(s_{t},\mathbf{x})$ are visit counts, and $c>0$ balances exploration and exploitation. The reward, defined as the total number of spheres is backpropagated to guide the MCTS algorithm toward configurations that maximize the kissing number. This process repeats until no feasible $\mathbf{x}$ remain, yielding the cosine set $\mathcal{C}_{1}=\{\cos\theta_{1},\dots,\cos\theta_{k}\}$ of allowed cosine values for possible kissing configuration. We note that a similar concept of the cosine set appeared in a previous note [?], but in our method it is used fundamentally differently, with the construction relying solely on cosine relationships.
Figure 2: Overview of PackingStar system. The workflow consists of three steps for $n$ dimensions: Step 1. New spheres tangent to n1n-1 existing spheres are iteratively solved from equations and added to the search space for simulation. Cosine frequencies are recorded until they converge to a discrete set to capture the geometric features. Step 2. The cosine set from Step 1 populates an initial matrix (red region) for the matrix completion game. Initialization can leverage geometric priors or start from scratch. Step 3. The first agent, Filler, fills the matrix with candidates, while the second agent, Corrector, removes previously added entries that are not good enough. After each round, the matrix is decomposed and reassembled for the next iteration based on geometric feature analysis.
Finding Kissing Numbers in two-player matrix completion game
Let $\mathbf{G}^{(m_{0})}\in\mathbb{R}^{m_{0}\times m_{0}}$ be the initial partial Gram matrix of the sphere center vectors. We formulate the matrix completion process as a two-player sequential game, corresponding to Step 2 and Step 3 in Figure 2. Each player is modeled as a learning agent, and the two agents are jointly optimized through game-theoretic reinforcement learning to cooperatively search for larger kissing numbers. This alternating filling-correcting process over $N$ rounds can be formulated as a Markov game, where the matrix $\mathbf{G}^{(m)}$ represents the state, Player 1’s action is the sampled entry $\mathbf{g}$ that extends the matrix, Player 2’s action is the selected index set $I$ for correcting entries, and the shared team reward is the number of rows of the final matrix.
Player 1 (Matrix Filler).
At each step $m\geq m_{0}$, Player 1 samples an action from policy $\pi_{1}$, $\mathbf{g}\sim\pi_{1}(\cdot\mid\mathbf{G}^{(m)})$, where $\mathbf{g}\in\mathcal{A}^{(m)}$ to extend the current matrix $\mathbf{G}^{(m)}\in\mathbb{R}^{m\times m}$:
$\mathbf{G}^{(m+1)}(\mathbf{g})=\begin{bmatrix}\mathbf{G}^{(m)}&\mathbf{g}\\ \mathbf{g}^{\top}&1\end{bmatrix}$. (4)
The constraints for candidate action set $\mathcal{A}^{(m)}$ depends on $m$, For $m<n$
$\mathcal{A}^{(m)}=\Big\{\mathbf{g}\in\mathcal{C}_{1}^{m}\;\Big|\;\mathbf{G}^{(m+1)}(\mathbf{G})\succeq 0,\;\mathrm{rank}(\mathbf{G}^{(m+1)}(\mathbf{g}))=m+1\Big\}.$ (5)
For $m\geq n$, split $\mathbf{g}$ into two parts:
$\mathbf{g}=\begin{bmatrix}\mathbf{g}^{(1)}\\ \mathbf{g}^{(2)}\end{bmatrix},\quad\mathbf{g}^{(1)}\in\mathbb{R}^{n},\ \mathbf{g}^{(2)}\in\mathbb{R}^{m-n}.$ (6)
Define the top-left block $\mathbf{G}^{(m)}_{1:n,\,1:n}\in\mathbb{R}^{n\times n}$ and its Cholesky factorization $\mathbf{G}^{(m)}_{1:n,\,1:n}=\mathbf{M}\mathbf{M}^{\top}$, with pseudo-inverse $\mathbf{M}^{+}$. Let
$\mathbf{G}^{\prime}=\mathbf{G}^{(m)}_{\,n+1:m,\,1:n}\in\mathbb{R}^{(m-n)\times n}.$ (7)
Compute the remaining part:
$\mathbf{g}^{(2)}={\mathbf{G}^{\prime}}(\mathbf{G}^{(m)}_{1:n,\,1:n})^{+}\mathbf{g}^{(1)}\,.$ (8)
Then the candidate action set is
$\mathcal{A}^{(m)}=\Big\{\mathbf{g}=[\mathbf{g}^{(1)};\mathbf{g}^{(2)}]\;\Big|\;\mathbf{g}^{(1)}\in\mathcal{C}_{1}^{n},\;\|\mathbf{M}^{+}\mathbf{g}^{(1)}\|_{2}=1,\;\mathbf{g}^{(2)}\in\mathcal{C}_{2}^{\,m-n}\Big\}.$ (9)
$\mathcal{C}_{2}$ is a predefined subset of allowed cosine values. $\mathcal{C}_{2}$ may coincide with $\mathcal{C}_{1}$, differ from it, or represent a continuous constraint set in different settings, e.g.,
$\mathcal{C}_{2}=\Big\{\mathbf{g}^{(2)}\in\mathbb{R}^{\,m-n}\;\Big|\;g^{(2)}_{i}\leq 0.5,\ i=1,\dots,m-n\Big\}.$ (10)
In specific cases, an additional constraint set $\mathcal{C}_{*}$ can be directly applied to the candidate action set,
$\mathcal{A}^{(m)}\leftarrow\mathcal{A}^{(m)}\cap\mathcal{C}_{*}$ (11)
Player 2 (Matrix Corrector).
Once Player 1 completes $M$ steps (or no feasible $\mathbf{g}$ remains), Player 2 samples an index set
$I\sim\pi_{2}(\cdot\mid\mathbf{G}^{(M)})\subseteq\{1,2,\dots,M\},$ (12)
and constructs the refined matrix
$\mathbf{G}^{*}=\mathbf{G}^{(M)}_{I,I}$. (13)
Game Dynamics and Cooperative Learning Process.
The policy $\pi_{2}$ is learned to identify and remove suboptimal entries filled by Player 1, so that in subsequent filling steps, Player 1 can expand the matrix further. Then the Player 1 resumes filling starting from $\mathbf{G}^{*}$, producing a new feasible set $\mathcal{A}^{(M-|I|)}$, and this alternation continues:
$\mathbf{G}^{(M)}\xrightarrow{\pi_{2}}\mathbf{G}^{*}\xrightarrow{\pi_{1}}\mathbf{G}^{(M^{\prime})}\xrightarrow{\pi_{2}}\cdots$ (14)
The alternating filling-correcting process continues until no further feasible entries exist for Player 1, and Player 2 cannot correct suboptimal entries. Let $\mathbf{G}^{(M)}$ denote the final completed matrix. The cooperative reinforcement learning objective of the PackingStar system is to maximize the shared team reward $R_{\text{team}}=\left\|\operatorname{diag}\bigl(\mathbf{G}^{(M)}\bigr)\right\|_{F}^{2}$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. The higher reward corresponds to a larger kissing number. Formally, the optimal joint policies $(\pi_{1}^{*},\pi_{2}^{*})$ are obtained by solving
$(\pi_{1}^{*},\pi_{2}^{*})=\arg\max_{\pi_{1},\pi_{2}}\;\mathbb{E}_{\mathbf{g}\sim\pi_{1},I\sim\pi_{2}}\biggl[\bigl\|\operatorname{diag}\bigl(\mathbf{G}^{(M)}(\{\mathbf{g}\},I)\bigr)\bigr\|_{F}^{2}\biggr],$ (15)
where the expectation accounts for the stochasticity of both policies, and $\mathbf{G}^{(M)}(\{\mathbf{g}\},I)$ denotes the matrix obtained after all filling and correcting steps under policies $\pi_{1}$ and $\pi_{2}$. In short, the policy of Player 1 is implemented as a tree-search learner, where node values are updated based on the team reward obtained after each episode. The policy of Player 2 is parameterized by a neural network and optimized via policy gradient. After each round, both the tree search statistics of Player 1 and the policy parameters of Player 2 are jointly updated under the same team reward, enabling the two players to mutually improve and construct larger kissing configurations. More implementation details will be provided in the supplementary materials of a future version.
Conclusion
Trained from scratch without any human-designed priors, PackingStar automatically discovers high-dimensional kissing configurations that outperform all known constructions and exhibit clear mathematical structure. The key idea is to reformulate the Kissing Number Problem as a two-player matrix completion game over the Gram-matrix domain, enabling efficient exploration of an extremely large and sparse space without explicit coordinate representations. Notably, PackingStar improves all known lower bounds from dimensions 25 to 31. In 13 dimensions, it uncovers a rational configuration with a larger kissing number; and in 14 and other dimensions, it identifies thousands of new configurations that provide a rich experimental basis for structural classification and better geometric analysis. A current limitation is that PackingStar cannot provide optimality proofs for these new configurations. We anticipate that the interaction between reinforcement learning and rigorous proof techniques will not only open new pathways for the Kissing Number Problem, but also extend to broader geometry problems.
Acknowledgments
We appreciate Ao Li’s assistance in program optimization and large-scale computing cluster deployment. We thank Mingzhi Wang and Juntao Dai for their assistance in coordinating GPU resources. We are grateful to Henry Cohn for his influential work on the Kissing Number Problem, which has been a source of inspiration for this research. We thank Chiyuan Wang and Xinmian Sun for the early discussions on this topic.
Author contributions
C.M. conceived and led the project. Z.T. established geometric metrics and analyzed the properties of the structures. C.M. developed the game-theoretic reinforcement learning method. M.L., Z.T., Z.M. and C.M. explored geometric relationships between structures of different dimensions. H.C. explored the training of geometric generative models for high-dimensional structures. C.M., H.C., and M.L. modeled the process of cosine feature simulation and designed the simulation algorithm. M.L., C.M. and Z.T. discovered the cosine feature of constructions via numerical simulation. C.M., H.C., Z.T., M.L., P.L., Z.M. and Y.Y. completed the formulation of matrix completion and designed the algorithm. Z.M. provided practical insights and metrics on the exploration of lattices and error-correcting codes, particularly in the construction of 31D, which provides core insights. C.M., P.L., M.L., Z.T. and H.C. developed the project code. C.M., Z.T., P.L. and H.C. conducted the experiments. C.M. and Y.Y. wrote the manuscript. Y.Q., Y.Y. and Y.C. provided computational resources and contributed to the discussions.
Competing interests
The authors declare no competing interests.

### v2

@@
 Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry and exponentially growing combinatorial complexity beyond 8 dimensions, which exceeds the complexity of Go game, limit the scalability of existing methods.
-Here we model this problem as a two-player matrix completion game and train the game-theoretic reinforcement learning system, PackingStar, to efficiently explore high-dimensional spaces.
+Here we model this problem as a two-player matrix completion game that can be fully parallelized at large scale, and train the game-theoretic reinforcement learning system, PackingStar, to efficiently explore high-dimensional spaces.
 The matrix entries represent pairwise cosines of sphere center vectors; one player fills entries while another corrects suboptimal ones, jointly maximizing the matrix size, corresponding to the kissing number.
@@
 PackingStar reproduces previous configurations and surpasses all human-known records from dimensions 25 to 31, with the configuration in 25 dimensions geometrically corresponding to the Leech lattice and suggesting possible optimality.
-It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions and discovers over 6000 new structures in 14 and other dimensions.
+It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions, discovers over 6000 new structures in 14 and other dimensions, and sets new records for generalized kissing configurations under various angular constraints.
 These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition and open new pathways for the Kissing Number Problem and broader geometry problems.
@@
 The sheer volume of these newly discovered arrangements provides rich opportunities for further investigation and may ultimately lead to configurations with even higher kissing numbers.
-Collectively, these results highlight AI’s ability to understand and reason about high-dimensional geometric structures, opening a new paradigm for studying the Kissing Number Problem and demonstrating the potential of AI to explore a broader class of challenging mathematical problems.
+In addition, PackingStar achieves significant new records for generalized kissing configurations under tighter angular constraints, including a 12-dimensional configuration with 81 spheres under the $1/4$ cosine constraint, as well as 14-dimensional and 17-dimensional configurations with 252 and 578 spheres under the $1/3$ constraint, respectively.
+All of these results have been independently validated and are now included in the standard databases for kissing numbers and spherical codes.
+In summary, these results highlight AI’s ability to understand and reason about high-dimensional geometric structures, opening a new paradigm for studying the Kissing Number Problem and demonstrating the potential of AI to explore a broader class of challenging mathematical problems.
 Key difference to AlphaEvolve
@@
 In contrast, PackingStar not only sets new bounds across higher dimensions but also discovers geometrically meaningful, analyzable configurations, providing concrete objects for further mathematical study [?].
+In summary, AlphaEvolve demonstrates that AI can “optimize” geometry, PackingStar shows that AI can truly “create” geometry.
+This represents the first genuinely systematic breakthrough of AI in high-dimensional discrete geometry and coding theory: it is not merely a numerical improvement, but a methodological advance, illustrating how AI can uncover new structures and insights in previously unexplored high-dimensional spaces.
 Results
 We evaluate the PackingStar system across multiple high-dimensional regimes and report new state-of-the-art kissing configurations.
-Our results fall into three categories: (i) new lower bounds in dimensions 25253131, (ii) improved rational constructions in $13$ dimensions, and (iii) diverse families of solutions in different dimensions.
+Our results fall into three categories: (i) new lower bounds in dimensions 25253131, (ii) improved rational constructions in $13$ dimensions, (iii) diverse families of solutions in different dimensions, and (iv) new records for generalized kissing configurations.
 These findings demonstrate that PackingStar can both recover classical lattice-based structures and uncover entirely new high-dimensional arrangements, substantially expanding the landscape of known kissing configurations.
@@
 previous $K_{\mathrm{prev}}(n)$ for $25\leq n\leq 31$.
-The asterisk $\bigstar$ indicates that under our new construction form, the theoretical lower bound in 31 dimensions should be 238350.
-The present version reports 238078 only due to computational and time limitations, rather than any intrinsic limitation of our method.
-The Future versions are expected to reach the theoretical bound.
-In the new construction form, bold items indicate the improved configuration forms discovered by PackingStar compared with previous forms.
+In the new construction form, bold items indicate the improved configuration forms discovered by PackingStar compared with previous forms [?].
    $n$    New construction form    $\boldsymbol{K_{\mathrm{new}}(n)}$    $K_{\mathrm{prev}}(n)$ [?]
@@
    30    $\boldsymbol{K(6)}+K(24)+2\sum_{i=1}^{24}|S_{i}|$    220440    219984
-   31    $\boldsymbol{K(7)}+K(24)+\boldsymbol{2\sum_{i=1}^{42}|S_{i}|}$       238078    232874
+   31    $\boldsymbol{K(7)}+K(24)+\boldsymbol{2\sum_{i=1}^{42}|S_{i}|}$    238350    232874
 In this setting, the first objective is to identify the largest possible subsets $S_{i}$.
-Each $S_{i}$ corresponds to a subset of shortest Leech lattice vectors whose pairwise cosine values are bounded by $1/4$; equivalently, in the PackingStar framework this corresponds to cosine sets $C_{1}=C_{2}=\{-1,\ 0,\ \pm 1/4\}$ together with the structural constraint that $C_{*}$ matches the Leech minimal-vector set.
+Each $S_{i}$ corresponds to a subset of shortest Leech lattice vectors whose pairwise cosine values are bounded by $1/4$; equivalently, in the PackingStar framework this corresponds to cosine sets $C_{1}=C_{2}=\{-1,\ 0,\ \pm 1/4\}$ together with the structural constraint that $C_{*}$ matches the Leech lattice minimal-vector set.
 Previous work identified subsets $|S_{i}|=488$ with no apparent geometric structure [?].
@@
 The resulting new lower bounds $K_{\mathrm{new}}(n)$ are summarized in Table 2.
-Compared with the previous best values $K_{\mathrm{prev}}(n)$, our constructions increase the kissing number by $8,38,68,152,1224,456,$ and $5204$ spheres in dimensions $25$ through $31$, respectively.
+Compared with the previous best values $K_{\mathrm{prev}}(n)$, our constructions increase the kissing number by $8,38,68,152,1224,456,$ and $5476$ spheres in dimensions $25$ through $31$, respectively.
 These improvements are also highlighted by the red markers in Fig.
@@
 Among the thousands of discovered variants, we consistently observe the presence of asymmetric structures with cosine set $C_{1}=C_{2}=\{-1,\ -3/4,\ 0,\ \pm 1/4,\ \pm 1/2\}$, similar to those identified in the 13-dimensional rational configuration.
+New records for generalized kissing configurations
+In this section, we apply PackingStar to explore generalized kissing configurations with maximal inner product at most $1/3$ and $1/4$.
+For geometric intuition, these cases may be viewed as double-sphere and triple-sphere kissing configurations, in which one asks how many additional unit spheres can be simultaneously tangent to two or three mutually tangent spheres, respectively.
+These configurations arise naturally as fundamental building blocks in sphere packings and play an important role in the study of the broader packing problem.
+The classical kissing number is denoted by $K(n)$, which corresponds to the case in which the pairwise angles between sphere centers have cosine at most $1/2$.
+More generally, for an angular constraint $1<\alpha<1$ with $\alpha\neq 1/2$, we use $K(n,\alpha)$ to denote maximum number of unit vectors such that the inner product between any two distinct vectors is at most $\alpha$.
+Table 4: New maximal values $K_{\mathrm{new}}(n,\alpha)$ for generalized kissing configurations with maximal inner product at most $\alpha$, obtained by PackingStar, compared with the previous best known values $K_{\mathrm{prev}}(n,\alpha)$.
+Bold entries indicate improved configurations discovered by PackingStar.
+$n$ $\alpha$ New cosine set $\boldsymbol{K_{\mathrm{new}}(n,\alpha)}$ $K_{\mathrm{prev}}(n,\alpha)$
+12 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 81 79
+14 $1/3$ $\{-1,\ 0,\ \pm\tfrac{1}{3}\}$; variants: $\{-1,\ -\tfrac{2}{3},\ 0,\ \pm\tfrac{1}{3}\}$ 252 240 [?]
+17 $1/3$ $\{-1,\ -\tfrac{7}{9},\ -\tfrac{5}{9},\ 0,\ \pm\tfrac{1}{9},\ \pm\tfrac{1}{3}\}$ 578 546 [?]
+As shown in Table 4, the new constructions from PackingStar yield new record for $K(n,\alpha)$ in several dimensions.
+In 12 dimensions under the $1/4$ cosine constraint, we achieve $K_{\mathrm{new}}(12,1/4)=81$, improving over the previous best of 79.
+Similarly, in 14 and 17 dimensions under the $1/3$ constraint, we obtain $K_{\mathrm{new}}(14,1/3)=252$ and $K_{\mathrm{new}}(17,1/3)=578$, surpassing the previously reported values of 240 and 546, respectively.
+Notably, the $K_{\mathrm{new}}(12,1/4)$ configuration contains no antipodal pairs; however, each of the 81 spheres has the same number of neighbors at each distance, forming a highly regular local structure.
+Such a combination of high local regularity without any antipodal pair appears to be rare in spherical code.
+We find several distinct variants of $K_{\mathrm{new}}(12,1/4)$ configuration, among which the most symmetric realization turns out to be the Kronecker product of the Schläfli spherical code with the unit equilateral triangle.
+This arrangement possesses a transitive symmetry group of order 311040.
+In dimension 17, PackingStar found a significant improvement: a configuration with 578 spheres, exceeding the previously known 546.
+All three geometrically non-isomorphic variants discovered share the same cosine set.
+Analysis of these variants reveals a remarkable structural property: the $K_{\mathrm{new}}(17,1/3)$ configuration can be interpreted as essentially two copies of the 16-dimensional 288-sphere configuration, which is known to be optimal under the cosine $1/4$ constraint, together with two additional points, i.e., $578=288\times 2+2$.
+This decomposition provides strong evidence that 578 is very likely optimal for the cosine $1/3$ constraint in dimension 17.
 Methods
@@
 Second, we formulate the search for maximal kissing numbers as a cooperative two-player matrix completion game, where one agent iteratively fills the Gram matrix and the other corrects suboptimal entries, jointly maximizing the final matrix size.
+This formulation allows all operations to be parallelized at large scale on NVIDIA GPUs using low-level mathematical operators in cuBLAS library, fully exploiting computational power to accelerate both the game simulation and model training.
 Identifying geometric features of the kissing configuration via simulation
@@
 Player 1 (Matrix Filler).
-At each step $m\geq m_{0}$, Player 1 samples an action from policy $\pi_{1}$, $\mathbf{g}\sim\pi_{1}(\cdot\mid\mathbf{G}^{(m)})$, where $\mathbf{g}\in\mathcal{A}^{(m)}$ to extend the current matrix $\mathbf{G}^{(m)}\in\mathbb{R}^{m\times m}$:
+For each step $m\geq m_{0}$, Player 1 samples an action according to policy $\pi_{1}$, $\mathbf{g}\sim\pi_{1}(\cdot\mid\mathbf{G}^{(m)})$, where $\mathbf{g}\in\mathcal{A}^{(m)}$ to extend the current matrix $\mathbf{G}^{(m)}\in\mathbb{R}^{m\times m}$:
 $\mathbf{G}^{(m+1)}(\mathbf{g})=\begin{bmatrix}\mathbf{G}^{(m)}&\mathbf{g}\\ \mathbf{g}^{\top}&1\end{bmatrix}$. (4)
@@
 For $m\geq n$, split $\mathbf{g}$ into two parts:
-$\mathbf{g}=\begin{bmatrix}\mathbf{g}^{(1)}\\ \mathbf{g}^{(2)}\end{bmatrix},\quad\mathbf{g}^{(1)}\in\mathbb{R}^{n},\ \mathbf{g}^{(2)}\in\mathbb{R}^{m-n}.$ (6)
+$\mathbf{g}=\begin{bmatrix}\mathbf{g}^{(1)}\\ \mathbf{g}^{(2)}\end{bmatrix},\quad\mathbf{g}^{(1)}\in\mathbb{R}^{n\times 1},\ \mathbf{g}^{(2)}\in\mathbb{R}^{(m-n)\times 1}.$ (6)
 Define the top-left block $\mathbf{G}^{(m)}_{1:n,\,1:n}\in\mathbb{R}^{n\times n}$ and its Cholesky factorization $\mathbf{G}^{(m)}_{1:n,\,1:n}=\mathbf{M}\mathbf{M}^{\top}$, with pseudo-inverse $\mathbf{M}^{+}$.
@@
 Compute the remaining part:
-$\mathbf{g}^{(2)}={\mathbf{G}^{\prime}}(\mathbf{G}^{(m)}_{1:n,\,1:n})^{+}\mathbf{g}^{(1)}\,.$ (8)
+$\mathbf{g}^{(2)}={\mathbf{G}^{\prime}}(\mathbf{G}^{(m)}_{1:n,\,1:n})^{+}\mathbf{g}^{(1)}\,$ (8)
 Then the candidate action set is
@@
 Once Player 1 completes $M$ steps (or no feasible $\mathbf{g}$ remains), Player 2 samples an index set
-$I\sim\pi_{2}(\cdot\mid\mathbf{G}^{(M)})\subseteq\{1,2,\dots,M\},$ (12)
-and constructs the refined matrix
+$I\sim\pi_{2}(\cdot\mid\mathbf{G}^{(M)})$ (12)
+where $\pi_{2}(\cdot\mid\mathbf{G}^{(M)})$ denotes a probability distribution on $\{1,2,\dots,M\}$, and constructs the refined matrix
 $\mathbf{G}^{*}=\mathbf{G}^{(M)}_{I,I}$. (13)
@@
 Then the Player 1 resumes filling starting from $\mathbf{G}^{*}$, producing a new feasible set $\mathcal{A}^{(M-|I|)}$, and this alternation continues:
-$\mathbf{G}^{(M)}\xrightarrow{\pi_{2}}\mathbf{G}^{*}\xrightarrow{\pi_{1}}\mathbf{G}^{(M^{\prime})}\xrightarrow{\pi_{2}}\cdots$ (14)
+$\mathbf{G}^{(M)}\xrightarrow{\pi_{2}}\mathbf{G}^{*}\xrightarrow{\pi_{1}}\mathbf{G}^{(M^{\prime})}\xrightarrow{\pi_{2}}\cdots.$ (14)
 The alternating filling-correcting process continues until no further feasible entries exist for Player 1, and Player 2 cannot correct suboptimal entries.
@@
 The key idea is to reformulate the Kissing Number Problem as a two-player matrix completion game over the Gram-matrix domain, enabling efficient exploration of an extremely large and sparse space without explicit coordinate representations.
-Notably, PackingStar improves all known lower bounds from dimensions 25 to 31.
-In 13 dimensions, it uncovers a rational configuration with a larger kissing number; and in 14 and other dimensions, it identifies thousands of new configurations that provide a rich experimental basis for structural classification and better geometric analysis.
+Notably, PackingStar improves all known lower bounds from dimensions 25 to 31 and sets new records for generalized kissing configurations under various angular constraints.
 A current limitation is that PackingStar cannot provide optimality proofs for these new configurations.
@@
 Acknowledgments
+We are grateful to Professor Henry Cohn for his influential work on the Kissing Number Problem, which has been a source of inspiration for this research.
+We are also deeply grateful for his guidance and valuable insights, particularly for helping us identify the new structures we discovered and providing constructive suggestions that greatly facilitated our progress in generalized kissing configurations.
 We appreciate Ao Li’s assistance in program optimization and large-scale computing cluster deployment.
 We thank Mingzhi Wang and Juntao Dai for their assistance in coordinating GPU resources.
-We are grateful to Henry Cohn for his influential work on the Kissing Number Problem, which has been a source of inspiration for this research.
 We thank Chiyuan Wang and Xinmian Sun for the early discussions on this topic.
@@
 conceived and led the project.
-Z.T.
+T.T.Z.
 established geometric metrics and analyzed the properties of the structures.
 C.M.
-developed the game-theoretic reinforcement learning method.
-M.L., Z.T., Z.M.
+established the matrix completion game and developed the game-theoretic reinforcement learning method.
+M.L., T.T.Z., Z.M.
 and C.M.
@@
 M.L., C.M.
-and Z.T.
+and T.T.Z.
 discovered the cosine feature of constructions via numerical simulation.
-C.M., H.C., Z.T., M.L., P.L., Z.M.
-and Y.Y.
+C.M., H.C., T.T.Z., M.L., P.L.
+and Z.M.
 completed the formulation of matrix completion and designed the algorithm.
 Z.M.
-provided practical insights and metrics on the exploration of lattices and error-correcting codes, particularly in the construction of 31D, which provides core insights.
-C.M., P.L., M.L., Z.T.
+provided practical insights and metrics for exploring lattices and error-correcting codes, particularly for the 31D construction.
+C.M., P.L., M.L., T.T.Z.
 and H.C.
 developed the project code.
-C.M., Z.T., P.L.
+C.M., T.T.Z., P.L.
 and H.C.
@@
 and Y.C.
-provided computational resources and contributed to the discussions.
+provided computational resources and improved the efficiency of AI systems in parallel computation.
 Competing interests

### v3

@@
 Since Isaac Newton first studied the Kissing Number Problem in 1694, determining the maximal number of non-overlapping spheres around a central sphere has remained a fundamental challenge.
-This problem represents the local analogue of Hilbert’s 18th problem on sphere packing, bridging geometry, number theory, and information theory.
-Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry and exponentially growing combinatorial complexity beyond 8 dimensions, which exceeds the complexity of Go game, limit the scalability of existing methods.
-Here we model this problem as a two-player matrix completion game that can be fully parallelized at large scale, and train the game-theoretic reinforcement learning system, PackingStar, to efficiently explore high-dimensional spaces.
-The matrix entries represent pairwise cosines of sphere center vectors; one player fills entries while another corrects suboptimal ones, jointly maximizing the matrix size, corresponding to the kissing number.
-This cooperative dynamics substantially improves sample quality, making the extremely large spaces tractable.
-PackingStar reproduces previous configurations and surpasses all human-known records from dimensions 25 to 31, with the configuration in 25 dimensions geometrically corresponding to the Leech lattice and suggesting possible optimality.
-It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions, discovers over 6000 new structures in 14 and other dimensions, and sets new records for generalized kissing configurations under various angular constraints.
-These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition and open new pathways for the Kissing Number Problem and broader geometry problems.
+This problem is the local analogue of Hilbert’s 18th problem, bridging geometry, number theory, and information theory.
+Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry, dimensional structure variability, and combinatorial explosion beyond Go limit the scalability and generality of existing methods.
+Here we model the problem as a two-player matrix completion game and train the reinforcement learning system, PackingStar, to play the games.
+The matrix entries represent pairwise cosines of sphere center vectors.
+One player fills entries while another corrects suboptimal ones to improve exploration quality, cooperatively maximizing the matrix size, corresponding to the kissing number.
+These matrices are decomposed into representative substructures, providing diverse bases and structural constraints that steer subsequent games and make extremely large spaces tractable.
+PackingStar surpasses records from dimensions 25 to 31 and sets new lower bounds for generalized kissing numbers under various angular constraints.
+It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions and discovers over 6000 new structures in other dimensions.
+Notably, some configurations challenge long-held antipodal paradigms, revealing algebraic correspondences with finite simple groups as well as geometric relationships across dimensions.
+Inspired by these patterns, humans devised further improved constructions.
+These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition via extreme-scale reinforcement learning and open new pathways for the Kissing Number Problem and broader geometry research.
 Introduction
@@
 In 1900, David Hilbert included the sphere packing problem in $n$-dimensional space as the 18th problem of his famous 23 problems [?].
-The Kissing Number Problem can be viewed as a local analogue of Hilbert’s 18th problem [?].
-Beyond geometry, it connects to number theory, group theory, and information theory [?,?,?].
+The Kissing Number Problem is the local analogue of Hilbert’s 18th problem [?].
+It connects to geometry, number theory, group theory, and information theory [?,?,?].
 Therefore, advancing this problem could foster progress across multiple fields.
@@
 In dimensions 25–31, studies conducted in 2011 and 2016 further improved the bounds by exploiting subsets of Leech lattice vectors combined with heuristic optimization [?,?].
-While these results mark significant milestones, they are limited by reliance on prior knowledge, and painstaking human design.
+While these results mark significant milestones, they are limited by reliance on prior knowledge, and specific human design.
 As dimensionality increases, the exponentially growing configuration space and the breakdown of full symmetry assumptions make existing approaches insufficient.
 These challenges highlight the need for novel paradigms to explore complex structures beyond human intuition.
-In this paper, we introduce PackingStar, a multi-agent reinforcement learning system that reformulates the Kissing Number Problem as a two-player matrix completion game.
-This approach scales efficiently across dimensions, mitigates the combinatorial explosion of high-dimensional search spaces, and overcomes the challenge of extreme data scarcity due to the limited human knowledge of high-dimensional constructions.
-In PackingStar, two cooperative agents learn to play the matrix completion game to efficiently explore irregular structures in high-dimensional spaces without human prior knowledge.
-The first agent focuses on incrementally filling matrix entries to construct larger and denser configurations.
-However, in the early stages of exploration, the candidate space for potential entries is exponentially large, and the agent only observes local matrix information, making erroneous placements inevitable.
-To address this problem, the second agent is introduced to learn how to identify and remove invalid or inconsistent entries.
-By leveraging a more global understanding of the matrix geometry and high-dimensional structure, this agent dynamically corrects the mistakes made by the first agent, enabling the joint system to converge toward larger configurations—corresponding to higher kissing numbers.
-The key insight to PackingStar’s success lies in two complementary design principles.
-First, the framework reformulates the kissing number construction as an equivalent matrix completion problem in the Gram matrix domain.
-This reformulation eliminates numerical instability and inefficiency inherent in high-dimensional coordinate representations, enabling all computations to operate directly on the Gram matrix in a fully parallelizable manner.
-The formulation seamlessly couples problem modeling with its learning methods, allowing efficient GPU acceleration and large-scale sampling for AI-driven discovery of kissing numbers.
-Second, PackingStar features the finely designed game-theoretic interaction between the two agents.
-By strategically performing delayed and selective corrections, the second agent imposes geometric constraints on the first agent’s expansive search, effectively regularizing exploration in an otherwise intractable space.
-This game dynamics not only improves the quality of sampled configurations and the efficiency of value estimation but also reduces the search complexity in exponentially large spaces, enabling scalable discovery of complex high-dimensional configurations that were previously inaccessible to existing methods.
+In this paper, we introduce PackingStar, a multi-agent reinforcement learning system that reformulates the Kissing Number Problem as a two-player matrix completion game in the Gram matrix domain.
+By operating directly on pairwise cosines rather than high-dimensional coordinates, this formulation eliminates numerical instability and computational inefficiency, enabling fully parallelizable GPU acceleration and large-scale sampling.
+Within PackingStar, the first agent incrementally fills matrix entries to construct larger and denser configurations.
+In early exploration stages, however, the candidate space is exponentially large and only local matrix information is observable, making erroneous placements inevitable.
+To address this challenge, the second agent performs delayed and selective corrections by leveraging a more global understanding of matrix, dynamically removing invalid or inconsistent entries and compressing the first agent’s search space.
+This interaction enables the joint system to converge toward larger configurations corresponding to higher kissing numbers.
+Meanwhile, the evolving matrices are iteratively decomposed into representative substructures that distill essential geometric patterns and impose structural constraints for subsequent games.
+Together, these mechanisms transform unconstrained combinatorial exploration into guided construction, substantially reducing reinforcement learning complexity in exponentially large spaces and enabling scalable discovery of high-dimensional configurations previously inaccessible to existing methods.
 Figure 1: AI breakthroughs in the Kissing Number Problem.
@@
 The blue triangle marks the 11D improvement achieved by AlphaEvolve (2025) [?].
-In 13D, PackingStar (red star) uncovers a new rational configuration with kissing number $K_{r}(13)=1146$, distinct from the current best non-rational lower bound $K(13)=1154$.
-For $25\leq n\leq 31$, the red stars indicate new lower bounds $K_{\mathrm{new}}(n)$ discovered by PackingStar, with the inset reporting the corresponding increments $\Delta K(n)=K_{\mathrm{new}}(n)-K_{\mathrm{prev}}(n)$ compared to the previous best $K_{\mathrm{prev}}(n)$.
+The red stars indicate new lower bounds $K_{\mathrm{new}}(n)$ discovered by PackingStar, with the inset reporting the corresponding increments $\Delta K(n)=K_{\mathrm{new}}(n)-K_{\mathrm{prev}}(n)$ compared to the previous best $K_{\mathrm{prev}}(n)$.
 PackingStar demonstrates remarkable scalability, successfully reproducing all known kissing number configurations in dimensions up to 31.
 Figure 1 illustrates some of PackingStar’s breakthrough discoveries on this problem.
-From dimensions 25 to 31, PackingStar surpasses the lower bounds reported in 2016 and also exceeds the meta-construction form introduced in 2011 [?,?], achieving what appears to be the optimal construction form along this path; the resulting construction form are summarized in Table 2.
+From dimensions 25 to 31, PackingStar surpasses the lower bounds reported in 2016 and also exceeds the construction form introduced in 2011 [?,?], achieving what appears to be the optimal construction form along this path; the resulting construction form are summarized in Table 2.
 Notably, the breakthrough configuration discovered in 25 dimensions exhibits a clear geometric pattern that corresponds closely to the substructure of Leech lattice.
@@
 The sheer volume of these newly discovered arrangements provides rich opportunities for further investigation and may ultimately lead to configurations with even higher kissing numbers.
-In addition, PackingStar achieves significant new records for generalized kissing configurations under tighter angular constraints, including a 12-dimensional configuration with 81 spheres under the $1/4$ cosine constraint, as well as 14-dimensional and 17-dimensional configurations with 252 and 578 spheres under the $1/3$ constraint, respectively.
+In addition, PackingStar achieves significant new records for generalized kissing configurations under tighter angular constraints, including 12-dimensional, 20-dimensional and 21-dimensional configurations with 81, 405 and 567 spheres under the $1/4$ cosine constraint, as well as 14-dimensional and 17-dimensional configurations with 252 and 578 spheres under the $1/3$ constraint, respectively.
+Notably, several of the newly discovered configurations fundamentally challenge long-held antipodal paradigms.
+Rather than conforming to symmetry-driven constructions centered around antipodal pairing, these configurations exhibit asymmetric yet highly regular geometric patterns.
+In multiple cases, they reveal deep algebraic correspondences with finite simple groups and previously unnoticed geometric relationships across dimensions.
+These structural insights have, in turn, inspired mathematicians to extract new and more general construction principles.
+Building on patterns uncovered by PackingStar, subsequent human-designed configurations have further improved existing lower bounds in several dimensions, including 22 dimensions, illustrating a productive feedback loop between AI-driven discovery and mathematical theory.
 All of these results have been independently validated and are now included in the standard databases for kissing numbers and spherical codes.
-In summary, these results highlight AI’s ability to understand and reason about high-dimensional geometric structures, opening a new paradigm for studying the Kissing Number Problem and demonstrating the potential of AI to explore a broader class of challenging mathematical problems.
-Key difference to AlphaEvolve
-The key difference between PackingStar and AlphaEvolve lies in the search space.
-AlphaEvolve optimizes explicit coordinates in $\mathbb{R}^{n}$, which is numerically unstable and scales exponentially with dimension; increasing computational resources cannot overcome this fundamental limitation.
-In contrast, PackingStar operates in the Gram-matrix domain with strict geometric constraints, enabling efficient large-scale sampling.
-This fundamental distinction explains why PackingStar scales to higher dimensions while AlphaEvolve cannot [?].
-Although AlphaEvolve improved the kissing number in 11 dimensions from 592 to 593, likely by adjusting the coordinates of the known 592-sphere configuration, the result lacks clear geometric structure and does not generalize to other dimensions.
-In contrast, PackingStar not only sets new bounds across higher dimensions but also discovers geometrically meaningful, analyzable configurations, providing concrete objects for further mathematical study [?].
-In summary, AlphaEvolve demonstrates that AI can “optimize” geometry, PackingStar shows that AI can truly “create” geometry.
-This represents the first genuinely systematic breakthrough of AI in high-dimensional discrete geometry and coding theory: it is not merely a numerical improvement, but a methodological advance, illustrating how AI can uncover new structures and insights in previously unexplored high-dimensional spaces.
+In summary, these results showcase AI’s power to understand high-dimensional geometry, reshape established mathematical intuitions, and drive progress on century-old geometry problems.
 Results
@@
 Based on this clear geometric interpretation, we conjecture that this subset $S_{i}$ may be optimal, and consequently, the 25-dimensional kissing configuration derived from this construction could also attain the maximum kissing number.
-The second objective is to identify improved meta construction forms that assemble these subsets into full configurations.
-Both objectives are jointly addressed within the unified PackingStar framework, which simultaneously increases the achievable size of $S_{i}$ and uncovers superior meta construction forms.
-Our findings show that PackingStar not only discovers larger subsets but also identifies a strictly better construction template than the meta-construction introduced in 2011; moreover, evidence from convergence and geometric consistency strongly suggests that the new template is optimal within this construction pathway.
+The second objective is to identify improved construction forms that assemble these subsets into full configurations.
+Both objectives are jointly addressed within the unified PackingStar framework, which simultaneously increases the achievable size of $S_{i}$ and uncovers superior construction forms.
+Our findings show that PackingStar not only discovers larger subsets but also identifies a better construction form; moreover, evidence from convergence and geometric consistency strongly suggests that the new template is optimal within this construction pathway.
 These advances yield especially pronounced improvements in dimensions 29 and 31.
@@
 Table 4 summarizes the cosine value among sphere pairs in the new configurations.
+New records for generalized kissing numbers
+In this section, we apply PackingStar to explore generalized kissing configurations with maximal inner product at most $1/3$ and $1/4$.
+For geometric intuition, these cases may be viewed as double-sphere and triple-sphere kissing configurations, in which one asks how many additional unit spheres can be simultaneously tangent to two or three mutually tangent spheres, respectively.
+These configurations arise naturally as fundamental building blocks in sphere packings and play an important role in the study of the broader packing problem.
+The classical kissing number is denoted by $K(n)$, which corresponds to the case in which the pairwise angles between sphere centers have cosine at most $1/2$.
+More generally, for an angular constraint $1<\alpha<1$ with $\alpha\neq 1/2$, we use $K(n,\alpha)$ to denote maximum number of unit vectors such that the inner product between any two distinct vectors is at most $\alpha$.
+Table 3: New maximal values $K_{\mathrm{new}}(n,\alpha)$ for generalized kissing configurations with maximal inner product at most $\alpha$, obtained by PackingStar, compared with the previous best known values $K_{\mathrm{prev}}(n,\alpha)$.
+Bold entries indicate improved configurations discovered by PackingStar, and $\star$ denotes human-improved constructions inspired by structures from PackingStar.
+$n$ $\alpha$ New cosine set $\boldsymbol{K_{\mathrm{new}}(n,\alpha)}$ $K_{\mathrm{prev}}(n,\alpha)$
+14 $1/3$ $\{-1,\ 0,\ \pm\tfrac{1}{3}\}$; variants: $\{-1,\ -\tfrac{2}{3},\ 0,\ \pm\tfrac{1}{3}\}$ 252 240 [?]
+17 $1/3$ $\{-1,\ -\tfrac{7}{9},\ -\tfrac{5}{9},\ 0,\ \pm\tfrac{1}{9},\ \pm\tfrac{1}{3}\}$ 578 546 [?]
+12 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 81 79 [?]
+20 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 405 378 [?]
+21 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 567 554 [?]
+22 $1/5$ $\{-1,\ -\tfrac{1}{5},\ \tfrac{1}{5}\}$ 352 338 [?]
+As shown in Table 3, the new constructions from PackingStar yield new record for $K(n,\alpha)$ in several dimensions.
+In 12, 20 and 21 dimensions under the $1/4$ constraint, we obtain $K_{\mathrm{new}}(12,1/4)=81$, $K_{\mathrm{new}}(20,1/4)=405$ and $K_{\mathrm{new}}(21,1/4)=567$, surpassing the previous records of 79, 378 and 554, respectively.
+We also report improvements under the $1/3$ constraint: in dimensions 14 and 17, we obtain $K_{\mathrm{new}}(14,1/3)=252$ and $K_{\mathrm{new}}(17,1/3)=578$, which beat the prior best known records of 240 and 546.
+Notably, the $K_{\mathrm{new}}(12,1/4)$ configuration contains no antipodal pairs; however, each of the 81 spheres has the same number of neighbors at each distance, forming a highly regular local structure.
+Such a combination of high local regularity without any antipodal pair appears to be rare in spherical code.
+We find several distinct variants of $K_{\mathrm{new}}(12,1/4)$ configuration, among which the most symmetric realization turns out to be the Kronecker product of the Schläfli spherical code with the unit equilateral triangle.
+This arrangement possesses a transitive symmetry group of order 311040.
+In dimension 17, PackingStar found a significant improvement: a configuration with 578 spheres, exceeding the previously known 546.
+All three geometrically non-isomorphic variants discovered share the same cosine set.
+Analysis of these variants reveals a remarkable structural property: the $K_{\mathrm{new}}(17,1/3)$ configuration can be interpreted as essentially two copies of the 16-dimensional 288-sphere configuration, which is known to be optimal under the cosine $1/4$ constraint, together with two additional points, i.e., $578=288\times 2+2$.
+This decomposition provides strong evidence that 578 is very likely optimal for the cosine $1/3$ constraint in dimension 17.
 Improved rational configurations in 13 dimensions
@@
 This reveals that even in 13 and higher dimensions, kissing configurations may possess special non-symmetric arrangements that were previously unexplored.
-Diverse constructions in different dimensions
-Beyond the specific breakthroughs in 13 and in dimensions 25 through 31, PackingStar uncovers a rich diversity of kissing configurations whose kissing number reaches current lower bound across multiple other dimensions.
-Among the thousands of discovered variants, we consistently observe the presence of asymmetric structures with cosine set $C_{1}=C_{2}=\{-1,\ -3/4,\ 0,\ \pm 1/4,\ \pm 1/2\}$, similar to those identified in the 13-dimensional rational configuration.
-New records for generalized kissing configurations
-In this section, we apply PackingStar to explore generalized kissing configurations with maximal inner product at most $1/3$ and $1/4$.
-For geometric intuition, these cases may be viewed as double-sphere and triple-sphere kissing configurations, in which one asks how many additional unit spheres can be simultaneously tangent to two or three mutually tangent spheres, respectively.
-These configurations arise naturally as fundamental building blocks in sphere packings and play an important role in the study of the broader packing problem.
-The classical kissing number is denoted by $K(n)$, which corresponds to the case in which the pairwise angles between sphere centers have cosine at most $1/2$.
-More generally, for an angular constraint $1<\alpha<1$ with $\alpha\neq 1/2$, we use $K(n,\alpha)$ to denote maximum number of unit vectors such that the inner product between any two distinct vectors is at most $\alpha$.
-Table 4: New maximal values $K_{\mathrm{new}}(n,\alpha)$ for generalized kissing configurations with maximal inner product at most $\alpha$, obtained by PackingStar, compared with the previous best known values $K_{\mathrm{prev}}(n,\alpha)$.
-Bold entries indicate improved configurations discovered by PackingStar.
-$n$ $\alpha$ New cosine set $\boldsymbol{K_{\mathrm{new}}(n,\alpha)}$ $K_{\mathrm{prev}}(n,\alpha)$
-12 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 81 79
-14 $1/3$ $\{-1,\ 0,\ \pm\tfrac{1}{3}\}$; variants: $\{-1,\ -\tfrac{2}{3},\ 0,\ \pm\tfrac{1}{3}\}$ 252 240 [?]
-17 $1/3$ $\{-1,\ -\tfrac{7}{9},\ -\tfrac{5}{9},\ 0,\ \pm\tfrac{1}{9},\ \pm\tfrac{1}{3}\}$ 578 546 [?]
-As shown in Table 4, the new constructions from PackingStar yield new record for $K(n,\alpha)$ in several dimensions.
-In 12 dimensions under the $1/4$ cosine constraint, we achieve $K_{\mathrm{new}}(12,1/4)=81$, improving over the previous best of 79.
-Similarly, in 14 and 17 dimensions under the $1/3$ constraint, we obtain $K_{\mathrm{new}}(14,1/3)=252$ and $K_{\mathrm{new}}(17,1/3)=578$, surpassing the previously reported values of 240 and 546, respectively.
-Notably, the $K_{\mathrm{new}}(12,1/4)$ configuration contains no antipodal pairs; however, each of the 81 spheres has the same number of neighbors at each distance, forming a highly regular local structure.
-Such a combination of high local regularity without any antipodal pair appears to be rare in spherical code.
-We find several distinct variants of $K_{\mathrm{new}}(12,1/4)$ configuration, among which the most symmetric realization turns out to be the Kronecker product of the Schläfli spherical code with the unit equilateral triangle.
-This arrangement possesses a transitive symmetry group of order 311040.
-In dimension 17, PackingStar found a significant improvement: a configuration with 578 spheres, exceeding the previously known 546.
-All three geometrically non-isomorphic variants discovered share the same cosine set.
-Analysis of these variants reveals a remarkable structural property: the $K_{\mathrm{new}}(17,1/3)$ configuration can be interpreted as essentially two copies of the 16-dimensional 288-sphere configuration, which is known to be optimal under the cosine $1/4$ constraint, together with two additional points, i.e., $578=288\times 2+2$.
-This decomposition provides strong evidence that 578 is very likely optimal for the cosine $1/3$ constraint in dimension 17.
 Methods
@@
 More implementation details will be provided in the supplementary materials of a future version.
+Matrix Decompositionn and Initialization for Subsequent Games.
+Once a final completed matrix $\mathbf{G}^{(M)}$ is obtained, it is iteratively decomposed into a collection of submatrices
+$\mathbf{G}^{(M)}_{J,J},\quad J\subset\{1,2,\dots,M\},$ (16)
+where each index set $J$ selects a subset of rows and columns from $\mathbf{G}^{(M)}$.
+These submatrices capture representative geometric patterns and serve as distilled structural bases.
+For subsequent filling games, each submatrix $\mathbf{G}^{(M)}_{J,J}$ is used as an initialization matrix, providing structured guidance and constraints for Player 1 while exploring new entries.
+Multiple such submatrices can be used to initialize parallel filling games, allowing the system to explore diverse regions of the high-dimensional search space simultaneously.
+Formally, for each submatrix $\mathbf{G}^{*}=\mathbf{G}^{(M)}_{J,J}$, Player 1 resumes filling according to policy $\pi_{1}$, and Player 2 applies corrections via $\pi_{2}$, continuing the alternating process.
 Conclusion
@@
 We are also deeply grateful for his guidance and valuable insights, particularly for helping us identify the new structures we discovered and providing constructive suggestions that greatly facilitated our progress in generalized kissing configurations.
-We appreciate Ao Li’s assistance in program optimization and large-scale computing cluster deployment.
-We thank Mingzhi Wang and Juntao Dai for their assistance in coordinating GPU resources.
+We appreciate the assistance of Ao Li and Xingmeng Zhang in program optimization and computing cluster deployment.
+We thank Mingzhi Wang and Juntao Dai in coordinating GPU resources.
 We thank Chiyuan Wang and Xinmian Sun for the early discussions on this topic.
@@
 T.T.Z.
-established geometric metrics and analyzed the properties of the structures.
+established geometric metrics and interpreted the properties of the structures.
 C.M.
@@
 explored the training of geometric generative models for high-dimensional structures.
-C.M., H.C., and M.L.
+C.M., H.C.
+and M.L.
 modeled the process of cosine feature simulation and designed the simulation algorithm.
@@
 discovered the cosine feature of constructions via numerical simulation.
-C.M., H.C., T.T.Z., M.L., P.L.
-and Z.M.
-completed the formulation of matrix completion and designed the algorithm.
+C.M., H.C., T.T.Z., M.L., P.L., Z.M.
+completed the formulation of matrix completion.
 Z.M.
-provided practical insights and metrics for exploring lattices and error-correcting codes, particularly for the 31D construction.
+provided practical insights and metrics for exploring lattices and error-correcting codes, particularly for the 31-dimensional construction.
 C.M., P.L., M.L., T.T.Z.
@@
 developed the project code.
+C.M.
+and P.L.
+optimized the efficiency of overall system.
 C.M., T.T.Z., P.L.
@@
 wrote the manuscript.
-Y.Q., Y.Y.
+Y.Q., Y.Y., and Y.C.
+provided computational resources and planned directions for improving the parallel efficiency.
+Y.Q.
 and Y.C.
-provided computational resources and improved the efficiency of AI systems in parallel computation.
+analyzed the computational time cost of the program, optimized the operator, and ensured the stable training.
 Competing interests
 The authors declare no competing interests.
+Supplementary Information
+Appendix A Key difference to AlphaEvolve
+The key difference between PackingStar and AlphaEvolve lies in the search space.
+AlphaEvolve optimizes explicit coordinates in $\mathbb{R}^{n}$, which is numerically unstable and scales exponentially with dimension; increasing computational resources cannot overcome this fundamental limitation.
+In contrast, PackingStar operates in the Gram-matrix domain with strict geometric constraints, combined with iterative decomposition into representative substructures and cooperative game dynamics that compress and distill high-dimensional search spaces.
+These techniques fundamentally render exploration of extremely large and complex spaces tractable, enabling efficient large-scale sampling.
+This fundamental distinction explains why PackingStar scales to higher dimensions while AlphaEvolve cannot [?].
+Although AlphaEvolve improved the kissing number in 11 dimensions from 592 to 593, likely by adjusting the coordinates of the known 592-sphere configuration, the result lacks clear geometric structure.
+Furthermore, AlphaEvolve relies on a specific 11-dimensional structure with 592 spheres, which is not jammed, making it difficult to transfer similar techniques to structures in other dimensions.
+In contrast, PackingStar not only sets new bounds across higher dimensions but also discovers geometrically meaningful, analyzable configurations, providing concrete objects for further mathematical study.
+In summary, AlphaEvolve demonstrates that AI can “optimize” geometry, PackingStar shows that AI can truly “create” geometry.
+This represents the first genuinely systematic breakthrough of AI in high-dimensional discrete geometry and coding theory: it is not merely a numerical improvement, but a methodological advance, illustrating how AI can uncover new structures and insights in previously unexplored high-dimensional spaces.
+Appendix B Diverse constructions in different dimensions
+Beyond the specific breakthroughs in 13 and in dimensions 25 through 31, PackingStar uncovers a rich diversity of kissing configurations whose kissing number reaches current lower bound across multiple other dimensions.
+Among the thousands of discovered variants, we consistently observe the presence of asymmetric structures with cosine set $C_{1}=C_{2}=\{-1,\ -3/4,\ 0,\ \pm 1/4,\ \pm 1/2\}$, similar to those identified in the 13-dimensional rational configuration.